OpenStudy (anonymous):

If f(x)=2−x^1/2 and g(x)=x^2−9, what is the domain of g(x)÷f(x)?

OpenStudy (owlcoffee):

Can you show me the result expression for \(\frac{ g(x) }{ f(x) }\)?

OpenStudy (anonymous):

Nope

OpenStudy (owlcoffee):

In the form I gave you, write down the expressions representing f(x) and g(x) respectively. That way you'll obtain the mathematical expression that represents the division of g(x) by f(x).

OpenStudy (anonymous):

Here's the problems and anwsers if i didnt type it right.

OpenStudy (owlcoffee):

I will not provide you with any answers, I will guide you throuh the solution of the problem.

OpenStudy (anonymous):

\[x^2-9/2-x^1/2\]

OpenStudy (anonymous):

I dont know how do use the equation thing but i know what you mean

OpenStudy (owlcoffee):

it is not "minus" but a division, meaning that, for instance: \[p \div q \iff \frac{ p }{ q }\] it's a way to express it.

OpenStudy (anonymous):

I dont understand

OpenStudy (owlcoffee):

if you are asked to express \(\frac{ g(x) }{ f(x) } \) and you know that the functions regarding g(x) and f(x) have the structures \(g(x)=x^2-9\) and \(f(x)=x ^{\frac{ 1 }{ 2 }}\) respectively. Then it's just a matter of plugging them in.

OpenStudy (anonymous):

I am plugging them in its not giving me an answer thats a choice.

OpenStudy (owlcoffee):

I am not telling you this is an answer, what I want you to do is to express the expression and then tell me what the domain of \((\frac{ g }{ f })(x)\). To put it more simple, in any rational function, what is the factor that can cause it to become undefined?

OpenStudy (anonymous):

I dont know.

OpenStudy (owlcoffee):

A more orienting question, what is the result of \(\frac{ 3 }{ 0 }\) ?

OpenStudy (anonymous):

0

OpenStudy (owlcoffee):

That's incorrect, try plugging it in a calculator.

OpenStudy (anonymous):

Well i dont know how to do it your not helping me at all just making me more confused

pooja195 (pooja195):

@EmiilyMxo appreciate the help you get if you dont understand say you dont understand rather than being rude. Tell the helper you dont understand you also need to make an effort to attempt to get the answer. Don't expect the answer to be given to you like a cookie. This is is openstudy not openanswers.

OpenStudy (anonymous):

I'm trying to get an answer thanks for the help anyway

pooja195 (pooja195):

Then I guess you're on your own. No one will want to help you if you are simply looking for answers. Can we try this again? And participate a bit more? ^.^

OpenStudy (anonymous):

Nope to late

pooja195 (pooja195):

Awww . Good luck in the future. :)

OpenStudy (owlcoffee):

I have already been issued by professors of online courses a couple of times for giving answers, and so did OS. And we are helpers, not answer givers; OS nor the helpers will contribute to you cheating in your assignments.

OpenStudy (anonymous):

Thanks for the "help" :))